OpenStudy (anonymous):

solve the quadratic equation? 3x^2+2x=0?

hartnn (hartnn):

0,-2/3.....need explanation?

OpenStudy (anonymous):

yess @hartnn show work!!

OpenStudy (anonymous):

Factor out x from that equation, x (3x+2) =0. So the factors are the places where the expression becomes zero.Can you guess it?

hartnn (hartnn):

x*(3x+2)=0 so x=0 and 3x+2=0....so x=0,x=-2/3

OpenStudy (anonymous):

you do notice thart that was a 3x^2 right??

OpenStudy (anonymous):

Yes of course. x*(3x+2) = 3x^2+2x .

OpenStudy (anonymous):

so the answer is x=0 x=-2/3

OpenStudy (anonymous):

Yes thats right.